# Benchmarking a Tunable Quantum Neural Network on Trapped-Ion and Superconducting Hardware

### v1

@@ Abstract
Ref. [5] presents theory and simulation of a quantum partially binarized multilayer perceptron i.e. the BQNN. See Fig. 1a. This network may be smoothly tuned between classical and quantum modes of operation. This network is partially binarized in the sense that, while each training weight can take on any real value, the activations of the hidden layers are constrained to be ±1$\pm$ 1± 1. For a classical network this is done by using the activation function
@@
where $\phi_{a}$ may be considered a quantum activation function. In the limit $a\rightarrow 0$, $\phi_{a}$ reduces to the classical activation function Eq. 1. In this case, each rotation angle is either 0 or ±$\pm\pi$. This means that, under ideal conditions, the rotation gates will only move the qubits between computational basis states, the behavior of the $a=0$ network is deterministic, and is equivalent to the standard classical neural net. Non-zero values of $a$ enable arbitrary angels of rotation and hence non-deterministic measurements. We identify $a$ as a parameter describing the “quantumness” of the network.
@@
The ion trap system is described in detail in Ref. [30]. It consists of ${}^{171}\mathrm{Yb}^{+}$ ions confined in a linear Paul trap. Each ion possesses two distinct types of quantum states: internal states and motional states. The internal states, used to encode qubits, correspond to hyperfine levels within the electronic ground-state manifold, ${}^{2}S_{1/2}$. Specifically, the qubit basis states |0|0$\rangle$| 0 ⟩ and |1|1$\rangle$| 1 ⟩ are defined as $|F,m_{F}\rangle=|0,0\rangle$ and |1,0|1,0$\rangle$| 1 , 0 ⟩ respectively. Here, $F$ indicates the total angular momentum including both the nuclear spin and total electronic angular momentum in these hyperfine-split states while $m_{F}$ is its projection onto the quantization axis. The motional states arise from the quantized collective oscillations of the ion chain within the harmonic potential formed by the radio-frequency and static electric fields of the Paul trap. These quantized vibrational modes serve as a shared quantum bus and play a central role in mediating interactions between qubits in trapped-ion quantum systems.
@@
Figure 2: (a) Validation rates on 55 randomly selected test images obtained from superconducting hardware and simulation for different values of $a$. For each image, ten shots were taken and the most common outcome was chosen as the predicted class. (b) An NY image that was misclassified at $a=0.0$ shows improved classification accuracy when the quantumness $a$ is increased to some optimal level on the three different platforms, i.e. superconducting, trapped-ion, and classical simulation. (c) For an NY and YY image, taken at either $a=0.0$ (classical) and $a=0.5$ (quantum), noise is controlled by the insertion of a number of single-qubit $UU^{\dagger}=I$ pairs.
@@
[…]

### v2

@@ Abstract
Ref. [5] presents theory and simulation of a quantum partially binarized multilayer perceptron i.e. the BQNN. See Fig. 1a. This network may be smoothly tuned between classical and quantum modes of operation. This network is partially binarized in the sense that, while each training weight can take on any real value, the activations of the hidden layers are constrained to be $\pm 1$. For a classical network this is done by using the activation function
@@
where $\phi_{a}$ may be considered a quantum activation function. In the limit $a\rightarrow 0$, $\phi_{a}$ reduces to the classical activation function Eq. 1. In this case, each rotation angle is either 0 or $\pm\pi$. This means that, under ideal conditions, the rotation gates will only move the qubits between computational basis states, the behavior of the $a=0$ network is deterministic, and is equivalent to the standard classical neural net. Non-zero values of $a$ enable arbitrary angels of rotation and hence non-deterministic measurements. We identify $a$ as a parameter describing the “quantumness” of the network.
@@
The ion trap system is described in detail in Ref. [30]. It consists of $^{171}\mathrm{Yb}^{+}$ ions confined in a linear Paul trap. Each ion possesses two distinct types of quantum states: internal states and motional states. The internal states, used to encode qubits, correspond to hyperfine levels within the electronic ground-state manifold, $^{2}S_{1/2}$. Specifically, the qubit basis states $|0\rangle$ and $|1\rangle$ are defined as $|F,m_{F}\rangle=|0,0\rangle$ and $|1,0\rangle$ respectively. Here, $F$ indicates the total angular momentum including both the nuclear spin and total electronic angular momentum in these hyperfine-split states while $m_{F}$ is its projection onto the quantization axis. The motional states arise from the quantized collective oscillations of the ion chain within the harmonic potential formed by the radio-frequency and static electric fields of the Paul trap. These quantized vibrational modes serve as a shared quantum bus and play a central role in mediating interactions between qubits in trapped-ion quantum systems.
@@
Figure 2: (a) Validation rates on 55 randomly selected test images obtained from superconducting hardware and simulation for different values of $a$. For each image, ten shots were taken and the most common outcome was chosen as the predicted class. (b) An NY image that was misclassified at $a=0.0$ shows improved classification accuracy when the quantumness $a$ is increased to some optimal level on the three different platforms, i.e. superconducting, trapped-ion, and classical simulation. (c) For an NY and $Y$ image, taken at either $a=0.0$ (classical) and $a=0.5$ (quantum), noise is controlled by the insertion of a number of single-qubit $UU^{\dagger}=I$ pairs.
@@
The final test uses native two-qubit gate pairs and results are depicted in Fig. 3. For the trapped-ion experiment it was found that after the introduction of a single pair of two-qubit gates, the per-image classification-rate shot up over $40\%$ before deteriorating to random 10% for five gate pairs. The superconducting platform was more resilient to this entanglement-noise and even showed a small improvement in performance with the addition of further pairs, a behavior not consistent with our intuition or simulations.